Find det(-2A) and compare it to det(A) for A=[-2 1 3][1 4 5][2 3 1]


1
Expert's answer
2021-06-16T09:31:07-0400

"det(A)=\begin{vmatrix}
 -2 & 1&3 \\
 1 & 4&5\\
2&3&1
\end{vmatrix}=2\cdot11+9-3\cdot5=16"


"det(-2A)=\begin{vmatrix}
 4 & -2&-6 \\
 -2 & -8&-10\\
-4&-6&-2
\end{vmatrix}=-4\cdot44-2\cdot36+6\cdot20=-128"


"det(2A)=-8det(A)"